40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Dec 7 18:47:36 2011, Jenne, Update, RF System, RAM Mon is now being demodulated 
    Reply  Thu Dec 8 00:04:50 2011, rana, Update, IOO, RAM Mon is now being demodulated 
       Reply  Thu Dec 8 14:47:28 2011, Jenne, Update, IOO, EOM aligned to minimize RAM 
    Reply  Thu Dec 8 00:45:13 2011, Koji, Update, RF System, 4ch demod is ready 
       Reply  Thu Dec 8 22:44:55 2011, Koji, Update, RF System, 4ch demod test result RF_DEMOD_TEST_111208.pdf
Message ID: 6092     Entry time: Thu Dec 8 22:44:55 2011     In reply to: 6086
Author: Koji 
Type: Update 
Category: RF System 
Subject: 4ch demod test result 

1) Linearity Test

LO input level was +10dBm. The LO freq was 11MHz and 55MHz for CH1 and CH2 respectively.
The IF frequency was fixed at 10kHz.

The amplitude of the RF input was swept from -50dBm to +15dBm.
Basically I and Q output of CH1 and CH2 was quite linear in this amplitude range.

2) Freqency Response

RF input was fixed at -20dBm and the IF frequency was swept from 1kHz to 1MHz.

The response was flat upto 100kHz, and have sensitivity upto 300kHz.

3) Output noise

Noise floor of the output is ~20nV/rtHz. All of the channels behave in the same way.
1/f start from 100Hz.

Attachment 1: RF_DEMOD_TEST_111208.pdf  103 kB  | Show | Show all
ELOG V3.1.3-